OpenStudy (amtran_bus):

If Kc = 0.405 at 40.°C and Kc = 0.575 at 90.°C, what is ΔH° for the reaction? RXN is simply x yields y and reverse

OpenStudy (frostbite):

Van Hoff equation mate.

OpenStudy (amtran_bus):

But do I need to rearrange it?

OpenStudy (frostbite):

You have to solve for \(\Delta H^{\circ}\) of cause. But the van't Hoff equation is given by: \[\Large \frac{ d \ln(K) }{ dT }=\frac{ \Delta H^{\circ} }{ RT^{2} }\] Solve the differential equation we get \[\Large \ln(\frac{ K_{2} }{ K_{1} })=\frac{ - \Delta H^{\circ} }{ R } \times \left( \frac{ 1 }{ T_{2} }-\frac{ 1 }{ T_{1} } \right)\] Solve for standard enthalpy and we are there.

OpenStudy (amtran_bus):

Can you help me solve for h? Can't you just use algebra instead of calculus?

OpenStudy (frostbite):

I don't see what the problem is. The last part is just algebra (solving for \(\Delta H\))?

OpenStudy (frostbite):

You don't have to solve the differential equation.. I've already done that for you.

OpenStudy (amtran_bus):

Thanks.

OpenStudy (frostbite):

Else let me fast show how I fast figure out how to solve the problem. We are interesting in some expression with \[\Large \frac{ \partial K }{ \partial T }\] We can relate the Gibbs free energy to the equilibrium constant from the known equation \[\Large \Delta G^{\circ}=-RT \ln(K)\]\[\Large \ln(K)=-\frac{ \Delta G^{\circ} }{ RT }\] Differentiation of ln K with respect to temperature \[\Large \frac{ d }{ dT }\ln(K)=-\frac{ 1 }{ R }\frac{ d }{ dT }\left( \frac{ \Delta G^{\circ} }{ T } \right)\] Using the Gibbs-Helmholtz expression we can evaluate the differential for Gibbs free energy. The Gibbs-Helmholtz expression is given by \[\Large \left( \frac{ \partial }{ \partial T } \frac{ \Delta G ^{\circ} }{ T } \right)_{p}=-\frac{ \Delta H ^{\circ} }{ T^2 }\] Substute we get the Van't Hoff equation \[\Large \Large \frac{ d \ln(K) }{ dT }=\frac{ \Delta H^{\circ} }{ RT^{2} }\] Now we need solve the equation. We choose to assume that \(\Delta H^{\circ}(T)\) is constant in the temperature interval [\(T_{1},T_{2}\)], we can then take \(\Delta H^{\circ}(T)\) out of the integral and treat it as a constant. Using separation of variables and integration in the respective intervals we get \[\Large \ln(T_{2})-\ln(T_{1})=\frac{ \Delta H^{\circ} }{ R }\int\limits_{T_{1}}^{T_{2}}T ^{-2} dT\] Solve the integral we get the solution I wrote \[\Large \ln(K_{2})-\ln(K_{1}) =\ \frac{ - \Delta H^{\circ} }{ R } \times \left( \frac{ 1 }{ T_{2} }-\frac{ 1 }{ T_{1} } \right)\] Solve for the standard enthalpy, \(\Delta H^{\circ}\) \[\Large \Delta H^{\circ}=-R\left( \frac{ \ln(K _{2})-\ln(K_{1}) }{\frac{ 1 }{ T_{2}}-\frac{ 1 }{ T_{1} } } \right)\]

OpenStudy (frostbite):

Sorry in the 7th expression I wrote \[\large \ln(T_{2})-\ln(T_{1})\] I meant \[\large \ln(K_{2})-\ln(K_{1})\] It should however not totally destroy the mathematical context.

OpenStudy (frostbite):

If you like to have fun with thermodynamics we can choose to change a bit in our approximations. In the previous method we chooses to assume that \[\Large \left( \frac{ \partial \Delta H^{\circ} }{ \partial T } \right)_{p}=\Delta C _{p}^{\circ}=0\] We can make a more exact calculation and this time assume that the change in standard heat capacity, \( \Delta C_{p}^{\circ}\) remains constant in the temperature interval [\(T_{1},T_{2}\)], so our approximation build on \[\Large \left( \frac{ \partial ^{2} \Delta H^{\circ} }{ \partial T^2 } \right)_{p}=\left( \frac{ \partial \Delta C _{p}^{\circ} }{ T } \right)_{p}=0\] Doing this we can solve the Van't Hoff equation to \[\ln(K_{2})-\ln(K_{1})=-\frac{ \Delta H^{\circ}(T_{1})- \Delta C_{p}^{\circ} T_{1} }{ R }\left( \frac{ 1 }{ T_{2} }-\frac{ 1 }{ T_{1} } \right)+\frac{ 1 }{ R } \Delta C_{p}^{\circ} \ln \left( \frac{ T_{2} }{ T_{1} } \right)\] This expression is more complex as we now take into account that the enthalpy is temperature dependent.

OpenStudy (frostbite):

Correction! \[\large \left( \frac{ \partial \Delta C_{p}^{\circ} }{ \partial T } \right)_{p}\] Forgot a partial sign the mathematical expression of the second possible approximation.